OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 3x + y - z = 2 2x + y + 2z = 8 2x - 3y + z = 21 (3, 4, 3) (-3, 4, 3) (3, -4, 3) (3, 4, -3)

OpenStudy (anonymous):

If we have : 8(3x + y - z = 2)=24x+8y-8z=16 and : 2( 2x + y + 2z = 8) = 4x+4y+8z=16 So can say : 24x+8y-8z=4x+4y+8z k ?:)

OpenStudy (anonymous):

@Futuristicmario

OpenStudy (texaschic101):

3x + y - z = 2 -->y = -3x + z + 2 2x + y + 2z = 8 2x + (-3x + z + 2) + 2z = 8 -x + 3z = 6 -x = -3x + 6 x = 3z - 6 y = -3x + z + 2 y = -3(3z - 6) + z + 2 y = -9z + 18 + z + 2 y = -8z + 20 2x - 3y + z = 21 2(3z - 6) - 3(-8z + 20) + z = 21 6z - 12 + 24z - 60 + z = 21 31z - 72 = 21 31z = 93 z = 3 x = 3z - 6 x = 3(3) - 6 x = 9 - 6 x = 3 y = -8z + 20 y = - 8(3) + 20 y = -24 + 20 y = - 4 check... 3x + y - z = 2 3(3) +(-4) - 3 = 2 9 - 4 - 3 = 2 9 - 7 = 2 2 = 2 (correct) check...2x - 3y + z = 21 2(3) - 3(-4) + 3 = 21 6 + 12 + 3 = 21 21 = 21 (correct) x = 3, y = -4, and z = 3